# Asymptotics for conformal inference

### v1

@@ 3 Asymptotics for conformal prediction
The aim here is to study the asymptotic properties of the processes $\mathrm{FCP}_{m}^{(n)}$ (5) and $\mathrm{FCP}_{m}^{w,(n)}$ (7). We study each process in two scenarios: exchangeable or distribution shift.
3.1 Main result
@@
where $F_{\tiny\mbox{cal}}^{-1}(u)=\inf\left\{x\in\mathbb{R}:F_{\tiny\mbox{cal}}(x)\geqslant u\right\}$ denotes the general inverse of $F_{\tiny\mbox{cal}}$. Formally, $I$ and $G$ correspond to the c.d.f. of the theoretical $p$-values $p^{(+\infty)}=1-F_{\tiny\mbox{cal}}(T)$ when $T\sim P_{\tiny\mbox{cal}}$ and $T\sim P_{\tiny\mbox{test}}$, respectively. We denote $G^{\prime}$ denotes the derivative of $G$ when it is defined.
@@
$\sqrt{\tau_{n,m}}\left(\mathrm{FCP}_{m}^{(n)}-I\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\text{ on $D[0,1]$,}$
@@
Theorem 3.1 is proved in Section 5.1.4. Note that Theorem 3.1 is true under slightly more general assumptions: namely, it is sufficient that all the scores are exchangeable without ties (Gazin et al.,, 2024). As a corollary, if $n/(n+m)\rightarrow\sigma^{2}\in(0,1]$, this result gives
$\sqrt{m}\left(\mathrm{FCP}_{m}^{(n)}-I\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma^{-1}\mathbb{U}.$
@@
Let us now consider the case where we have potentially a covariate shift, that is, $F_{\tiny\mbox{cal}}\neq F_{\tiny\mbox{test}}$. For this, we consider the following additional assumptions [^3] Assumption 3 can be slightly relaxed, by only assuming that $F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}$ is continuously differentiable on $(0,1)$. on $F_{\tiny\mbox{cal}}$ and $F_{\tiny\mbox{test}}$.
@@
$F_{\tiny\mbox{cal}}$ is increasing on its support $(a,b)\subset\mathbb{R}$ with $-\infty\leqslant a<b\leqslant+\infty$ and is continuously differentiable. In addition, $F_{\tiny\mbox{test}}$ is continuously differentiable.
@@
$\sqrt{\tau_{n,m}}\left(\mathrm{FCP}_{m}^{(n)}-G\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}\circ G+\sqrt{1-\sigma^{2}}G^{\prime}\mathbb{V}\text{ on $D(0,1)$},$
@@
Theorem 3.2 is proved in Section 5.1.3. When $F_{\tiny\mbox{test}}=F_{\tiny\mbox{cal}}$ it recovers Theorem 3.1 (under stronger assumptions). Otherwise, it shows that the $\mathrm{FCP}$ converges to $G\neq I$ which entails that $\mathrm{FCP}_{m}^{(n)}(\alpha)$ is not provided to converge to $\alpha$ anymore.
@@ 3.2 Weighted case
which corresponds to the c.d.f. of the distribution induced by weight function $w$ on the distribution $P_{\tiny\mbox{cal}}$. Note that when choosing the oracle weights (8), the latter is simply the c.d.f. of $P_{\tiny\mbox{test}}$. Second, let
@@
the c.d.f. of the theoretical $p$-values $p^{w,(+\infty)}=1-{F^{w}_{{\tiny\mbox{cal}}}}(T)$ with $T\sim P_{\tiny\mbox{test}}$, and we denote $(G^{w})^{\prime}$ its derivatives when it exists. Finally, we introduce quantities involved in the variance:
${\rho^{w}}$ $\coloneqq\frac{\left(\int_{\mathbb{R}}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)\right)^{1/2}}{\int_{\mathbb{R}}w(u)\operatorname{d}P_{\tiny\mbox{cal}}(u)}\geqslant 1;$ (16)
${V}^{w}_{{\tiny\mbox{cal}}}(t)$ $\coloneqq\frac{\int_{-\infty}^{t}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)}{\int_{\mathbb{R}}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)},\ t\in\mathbb{R};$ (17)
$I^{w}(t)$ $\coloneqq 1-{V}^{w}_{{\tiny\mbox{cal}}}\circ{({F^{w}_{{\tiny\mbox{cal}}}})^{-1}}(1-t),\ t\in(0,1).$ (18)
@@
The weight function $w:\mathbb{R}\cup\{\infty\}\rightarrow\mathbb{R}^{+}$ is uniformly bounded by a constant and is measurable. Moreover, ${F^{w}_{{\tiny\mbox{cal}}}}$ is increasing on its support $(a^{w},b^{w})\subset\mathbb{R}$ with $-\infty\leqslant a^{w}<b^{w}\leqslant+\infty$ and is continuously differentiable.
Typically, Assumption 4 holds if $\left\{t\in\mathbb{R},\ w(t)>0\right\}$ is a (possibly infinite) interval of $\mathbb{R}$ on which $w$ is bounded and continuous and $F_{\tiny\mbox{cal}}$ satisfies Assumption 3.
@@
$\sqrt{\tau_{n,m}}\left(\mathrm{FCP}_{m}^{w,(n)}-G^{w}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}\circ G^{w}+\sqrt{1-\sigma^{2}}{\rho^{w}}(G^{w})^{\prime}\left(\mathbb{V}\left(I^{w}\right)+\left[I-I^{w}\right]N\right)\text{ on $D(0,1)$},$
where $\mathbb{U}$, $\mathbb{V}$ are two independent standard Brownian bridges, ${\rho^{w}}$ is defined by (16), $I^{w}$ is defined by (18), $N$ is an independent standard Gaussian random variable and $(G^{w})^{\prime}$ denotes the derivative of $G^{w}$.
Theorem 3.3 is proved in Section 5.1.5. Note that Theorem 3.3 recovers Theorem 3.2 when $w\equiv 1$. The asymptotic expectation of $\mathrm{FCP}_{m}^{w,(n)}(\alpha)$ is $G^{w}(\alpha)$ which is different from $\alpha$ in general, if $w$ is chosen arbitrarily. Interestingly, when choosing the oracle weight function $w^{*}$ given by (8), we have $G^{w^{*}}=I$ and the convergence in Theorem 3.3 reads as follows:
@@
$\sqrt{\tau_{n,m}}\left(\mathrm{FCP}_{m}^{w^{*},(n)}-I\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}+\sqrt{1-\sigma^{2}}{\rho^{w^{*}}}\left(\mathbb{V}\left(I^{w^{*}}\right)+\left[I-I^{w^{*}}\right]N\right)\text{ on $D(0,1)$},$
@@
This result means that we can recover a result close to Theorem 3.1 in case of a distribution shift if the oracle weight function is used in the prediction sets.
Finally, the covariance terms of Theorems 3.3 and 3.1 can be used to build confidence intervals for the FCP, as illustrated in Figure 2.
@@ 4.1 Additional notation and assumptions
We denote, for all $m\geqslant 1$, $m_{0}(m)=|{\mathcal{H}}_{0}\cap\llbracket m\rrbracket|$ and $\pi_{0}(m)=m_{0}(m)/m$ the number and the proportion of nulls (i.e. of non-novelties) among $m$ tested points, respectively. We introduce the following quantities:
$G_{0}(t)$ $=1-F_{0}\circ F_{\tiny\mbox{cal}}^{-1}(1-t),\ t\in(0,1);$ (19)
@@
that correspond to the (limiting) c.d.f. of the conformal $p$-values under the null and under the test sample mixture, respectively. For some weight function $w$, we also define their weighted counterparts:
$G_{0}^{w}(t)$ $=1-F_{0}\circ{\left({F^{w}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-t),\ t\in(0,1),$ (21)
@@
The two following assumptions are classical when studying the asymptotic of multiple testing procedures (Genovese and Wasserman,, 2002).
@@
As shown in Chi, (2007); Neuvial, (2008) in the independent case, there is a critical value for the BH procedure given by $\alpha^{*}=[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$; the central limit theorem for the FDP/TDP of BH procedure at level $\alpha$ can only be obtained if $\alpha>\alpha^{*}$ (for $\alpha<\alpha^{*}$, the BH procedure at level $\alpha$ has asymptotically no power). We will show below that this quantity plays a similar role in the (dependent) conformal setting and prove results only if $\alpha>\alpha^{*}$.
Finally, as in Neuvial, (2008), we consider also the following thresholds
$\mathcal{T}_{\alpha}$ $=\sup\left\{t\in(0,1),\ G_{{\tiny\mbox{mixt}}}(t)\geqslant\frac{t}{\alpha}\right\};$ (23)
$\mathcal{T}^{w}_{\alpha}$ $=\sup\left\{t\in(0,1),\ G_{{\tiny\mbox{mixt}}}^{w}(t)\geqslant\frac{t}{\alpha}\right\},$ (24)

### v3

@@ 3 Asymptotics for conformal prediction
The aim here is to study the asymptotic properties of the processes $\mathrm{FCP}_{m}^{(n)}$ (6) and $\mathrm{FCP}_{m}^{w,(n)}$ (8). We study each process in both exchangeable or distribution shift framework.
3.1 Main results
@@
where $F_{\tiny\mbox{cal}}^{-1}(u)=\inf\mathopen{}\mathclose{{\left\{x\in\mathbb{R}:F_{\tiny\mbox{cal}}(x)\geqslant u}}\right\}$ denotes the general inverse of $F_{\tiny\mbox{cal}}$. Formally, $I$ and $G$ correspond to the c.d.f. of the theoretical $p$-values $p^{(+\infty)}=1-F_{\tiny\mbox{cal}}(T)$ when $T\sim P_{\tiny\mbox{cal}}$ and $T\sim P_{\tiny\mbox{test}}$, respectively. We denote $G^{\prime}$ the derivative of $G$ provided that it is well defined.
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\text{ on $D[0,1]$,}$
@@
Remark 3.1.
Theorem 3.1 is true under slightly more general assumptions: namely, it is sufficient that all the scores are exchangeable and without ties almost surely (see Theorem 5.2 for a formal statement). Using the latter is useful since non i.i.d. scores arise naturally in various setting, like in link prediction for graph (Blanchard et al., 2024) or when using adaptive scores (Gazin, Blanchard and Roquain, 2024).
Theorem 3.1 is proved in Section 5.1.3. As a corollary, if $n/(n+m)\rightarrow\sigma^{2}\in(0,1]$, this gives
$\sqrt{m}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma^{-1}\mathbb{U}.$
@@
Theorem 3.1 (or more precisely its extension Theorem 5.2 when the scores are exchangeable and without ties almost surely) can be formulated as a result on the asymptotic properties of the empirical c.d.f. of the random vectors $(q_{1},\cdots,q_{m})$ with the distribution $P_{n,m}$ defined in Gazin, Blanchard and Roquain (2024). More formally, let us assume that we are doing $m$ sucessive draws in a Pólya urn model with initially $1$ ball of each of the $n+1$ colors labelled $\mathopen{}\mathclose{{\left\{i/(n+1),\ i\in\llbracket n+1\rrbracket}}\right\}$. We denote $(q_{i}^{(n)})_{i\in\llbracket m\rrbracket}$ the colors of the $m$ successive draws. Then Theorem 3.1 yields,
$\Bigg(\sqrt{\tau_{n,m}}\bigg[m^{-1}\sum_{i\in\llbracket m\rrbracket}{\mathds{1}_{q_{i}^{(n)}\leqslant t}}-t\bigg]\Bigg)_{t\in[0,1]}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\ \text{on }D[0,1],$
with $\mathbb{U}$ being a standard Brownian Bridge. Since obtaining this Pólya urn consists on a first sampling of $n+m$ i.i.d. real random variables with a continuous c.d.f., and then sample without replacement $n$ points to compute the ranks of each of the $m$ other points, Theorem 3.1 can also be used to obtain the asymptotic of this kind of sampling without replacement procedure. Another interesting application is the convergence of the uniform distribution over the set of discrete random measure on the set $\mathopen{}\mathclose{{\left\{i/(n+1),\ i\in\llbracket n+1\rrbracket}}\right\}$ with $m$ atoms to the uniform distribution on $(0,1)$ whenever $n\wedge m\rightarrow+\infty$. Formally, let $Q_{m}^{(n)}$ be a random probability measure drawn uniformly from the set of discrete random measure on the set $\mathopen{}\mathclose{{\left\{i/(n+1),\ i\in\llbracket n+1\rrbracket}}\right\}$ with $m$ atoms. Denoting for all $t\in[0,1]$, $F_{m}^{(n)}(t)=Q_{m}^{(n)}([0,t])$ its c.d.f. Theorem 3.1 yields the following convergence in distribution
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(F_{m}^{(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\mathbb{U}\ \text{on $D[0,1]$},$
with $\mathbb{U}$ being a standard Brownian Bridge. This gives that the random measure $(Q_{m}^{(n)})_{m,n}$ converges in probability to the uniform distribution on $(0,1)$ for the Kolmogorov Smirnov divergence, with an asymptotic quantification of its fluctuations. All those convergences are direct consequences a direct corollary of Theorem 3.1 (and Theorem A in the supplementary material of Gazin, Blanchard and Roquain (2024) giving us various properties of $P_{n,m}$). However, establishing such links between conformal inference and Pólya urn theory in the event of a distribution shift is neither straightforward nor direct.
Let us now consider the case where we have potentially a distribution shift, that is, $F_{\tiny\mbox{cal}}\neq F_{\tiny\mbox{test}}$. For this, we consider the following additional assumptions on $F_{\tiny\mbox{cal}}$ and $F_{\tiny\mbox{test}}$.
@@
$F_{\tiny\mbox{cal}}$ is increasing on its support $(a,b)\subset\mathbb{R}$ with $\infty\leqslant a<b\leqslant+\infty$ and is continuously differentiable. In addition, $F_{\tiny\mbox{test}}$ is continuously differentiable.
Remark 3.2.
Assumptions on $F_{\tiny\mbox{cal}}$ and $F_{\tiny\mbox{test}}$ (Assumption 3 and the continuity assumption of Assumption 1) can be slightly relaxed by rather imposing a condition on the marginal distribution $F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}-}^{-1}$ of the theoretical $p$-value (5): Section G is devoted to the formal statement of those relaxed theorems.
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{(n)}-G}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}\circ G+\sqrt{1-\sigma^{2}}G^{\prime}\mathbb{V}\text{ on $D(0,1)$},$
@@
Theorem 3.2 is proved in Section 5.1.4. When $F_{\tiny\mbox{test}}=F_{\tiny\mbox{cal}}$ it recovers Theorem 3.1 (under the additional convergence assumption of $n/(n+m)$) since if $\mathbb{U}$ and $\mathbb{V}$ are two independent standard Brownian bridges and $\sigma\in[0,1]$, then $\sigma\mathbb{U}+\sqrt{1-\sigma^{2}}\mathbb{V}$ is also a standard Brownian bridge. By contrast, when $F_{\tiny\mbox{test}}\neq F_{\tiny\mbox{cal}}$, Theorem 3.2 quantifies the effect of the distribution shift. In particular, the limit of $\mathrm{FCP}_{m}^{(n)}$ is $G$ instead of simply $I$, thus $\mathrm{FCP}_{m}^{(n)}(\alpha)$ is not provided to converge to $\alpha$ anymore.
The conclusions of Theorems 3.1 and 3.2 involve a double asymptotic in $n$ and $m$ for which $n/(n+m)\rightarrow\sigma^{2}\in[0,1]$. The different regimes are summarized in Table 1. Whenever $\sigma^{2}$ is neither 0 nor $1$ this is equivalent to have $m\sim_{\tau_{n,m}\rightarrow\infty}(\sigma^{-2}-1)n$ and thus $m$ and $n$ are asymptotically proportional. For $\sigma^{2}=0$ (‘under-calibrated’ regime), the convergence rate is slower, of order $\sqrt{n}\ll\sqrt{m}$, for instance $m^{1/4}$ for $n=\sqrt{m}$, or $\sqrt{m(\log(m))^{-1}}$ for $n=m(\log(m))^{-1}$. This corresponds to the case where the conformal $p$-values are ‘strongly’ dependent. On the other hand, if $\sigma^{2}=1$ (‘over-calibrated’ regime), the rate $\sqrt{\tau_{n,m}}$ is equivalent to the classical rate in case of independence $\sqrt{m}$, with the same asymptotic variance. In that case, the convergence result is the same as for the independent theoretical $p$-values (5). This observation still apply in the weighted conformal prediction case.
Regime of convergence Rate of convergence: $\sqrt{\tau_{n,m}}$ Variance term
Conformal Under-calibrated: $m/n\rightarrow+\infty$ $\sqrt{n}$ $\sigma^{2}=0$
Proportional: $n\sim\gamma m,\ \gamma>0$ $\sqrt{\frac{\gamma}{\gamma+1}m}$ $\sigma^{2}={\frac{\gamma}{\gamma+1}}\in(0,1)$
Over-calibrated: $n\rightarrow 0 m\sqrt{{m}}$ $\sigma^{2}=1$
Independent Theoretical $p$-values (as $n=+\infty$) $\sqrt{m}$ as $\sigma^{2}=1$
Table 1: Three different regimes of convergence: $\sigma^{2}=0$, $\sigma^{2}\in(0,1)$ and $\sigma^{2}=1$ linked with the ratio between the number of calibration points and test points. The case of theoretical independent $p$-values is presented for comparison.
@@ 3.2 Weighted case
which corresponds to the c.d.f. of the distribution induced by weight function $w$ on the distribution $P_{\tiny\mbox{cal}}$ (i.e. proportional to $w(u)\operatorname{d}P_{\tiny\mbox{cal}}(u)$). Note that when choosing the oracle weight function (9), the latter is simply the c.d.f. of $P_{\tiny\mbox{test}}$. Second, we let
@@
the c.d.f. of the theoretical $p$-values $p^{w,(+\infty)}=1-{F^{w}_{{\tiny\mbox{cal}}}}(T)$ with $T\sim P_{\tiny\mbox{test}}$, and we denote $(G^{w})^{\prime}$ its derivative when it exists. Finally, we introduce quantities involved in the asymptotic variance:
${\rho^{w}}$ $\coloneqq\frac{\mathopen{}\mathclose{{\left(\int_{\mathbb{R}}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)}}\right)^{1/2}}{\int_{\mathbb{R}}w(u)\operatorname{d}P_{\tiny\mbox{cal}}(u)}\geqslant 1;$ (17)
$I^{w}(t)$ $\coloneqq 1-{V}^{w}_{{\tiny\mbox{cal}}}\circ{({F^{w}_{{\tiny\mbox{cal}}}})^{-1}}(1-t),\ t\in(0,1),$ (18)
with
${V}^{w}_{{\tiny\mbox{cal}}}(t)\coloneqq\frac{\int_{-\infty}^{t}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)}{\int_{\mathbb{R}}w(u)^{2}\operatorname{d}P_{\tiny\mbox{cal}}(u)},\ t\in\mathbb{R}.$ (19)
@@
The weight function $w:\mathbb{R}\cup\{\infty\}\rightarrow\mathbb{R}^{+}$ is uniformly bounded by a constant and is measurable. Moreover, $F^{w}_{{\tiny\mbox{cal}}}$} is increasing on its support $(a^{w},b^{w})\subset\mathbb{R}$ with $\infty\leqslant a^{w}<b^{w}\leqslant+\infty$ and is continuously differentiable.
Typically, Assumption 4 holds if $\mathopen{}\mathclose{{\left\{t\in\mathbb{R},\ w(t)>0}}\right\}$ is a possibly infinite interval of $\mathbb{R}$ on which $w$ is bounded and continuous, and $F_{\tiny\mbox{cal}}$ satisfies Assumption 3.
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{w,(n)}-G^{w}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}\circ G^{w}+\sqrt{1-\sigma^{2}}{\rho^{w}}(G^{w})^{\prime}\mathopen{}\mathclose{{\left(\mathbb{V}\circ{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right)\text{ on $D(0,1)$},$
where $\mathbb{U}$, $\mathbb{V}$ are two independent standard Brownian bridges, NN is an independent standard Gaussian random variable, $\rho^{w}$} is defined by (17) and $I^{w}$ is defined by (18).
Theorem 3.3 is proved in Section 5.1.5. Note that it recovers Theorem 3.2 when $w\equiv 1$. The asymptotic expectation of $\mathrm{FCP}_{m}^{w,(n)}(\alpha)$ is $G^{w}(\alpha)$ which is different from $\alpha$ for a general $w$. Interestingly, when choosing the oracle weight function $w^{*}$ given by (9), we have $G^{w^{*}}=I$ and the convergence in Theorem 3.3 reads as follows:
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FCP}_{m}^{w^{*},(n)}-I}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\sigma\mathbb{U}+\sqrt{1-\sigma^{2}}{\rho^{w^{*}}}\mathopen{}\mathclose{{\left(\mathbb{V}\circ{I^{w^{*}}}+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)\text{ on $D(0,1)$},$
@@
Theorem 3.4 shows that, in the case of a perfect training of the weight function, one can recover the same asymptotic behavior as in case of no distribution shift. However, note that the asymptotic variance is no longer universal and depends on the distributions proportionnal to $w\operatorname{d}P_{\tiny\mbox{cal}}$ and $w^{2}\operatorname{d}P_{\tiny\mbox{cal}}$. By contrast, Theorem 3.3 quantifies the impact of the distribution shift when the weight function is mis-specified, with a limit $G^{w}$ and not $I$. Comparing this to Theorem 3.2, Theorem 3.3 indicates that, for a level $\alpha$, the FDP behavior is (asymptotically) more favorable if we consider a weight function such that $G^{w}(\alpha)$ is closer to $\alpha$ than $G(\alpha)$. This phenomenon is illustrated in Figure 2 in a particular setting.
Finally, the covariance terms of Theorems 3.3 and 3.2 can be used to build asymptotic confidence intervals for the FCP, as illustrated in Figure 2.
@@ 4.1 Additional notation and assumptions
[…]
$G(t)$ $=1-F_{\tiny\mbox{test}}\circ F_{\tiny\mbox{cal}}^{-1}(1-t),\ t\in(0,1),$
$G_{0}(t)$ $=1-F_{0}\circ F_{\tiny\mbox{cal}}^{-1}(1-t),\ t\in(0,1),$ (20)
@@
that correspond to the (limiting) c.d.f. of the theoretical $p$-values (5) under the alternative, the null and under the test sample mixture, respectively. For a given weight function $w$, we also define their weighted counterparts and recall the definition of $G^{w}$ (16):
$G^{w}(t)$ $=1-F_{\tiny\mbox{test}}\circ({F^{w}_{{\tiny\mbox{cal}}}})^{-1}(1-t),\ t\in(0,1),$
$G_{0}^{w}(t)$ $=1-F_{0}\circ{\mathopen{}\mathclose{{\left({F^{w}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-t),\ t\in(0,1),$ (22)
@@
Let us introduce the two following assumptions:
@@
Since $G_{{\tiny\mbox{mixt}}}$ can be seen as the cdf of the theoretical $p$-values (5) in the test sample, Assumption 5 is classical in the multiple testing literature (Genovese and Wasserman, 2002) (see Section G for more precision), and Assumption 6 corresponds to its weighted version.
As shown in Chi (2007); Neuvial (2008) in the independent case, there is a critical value for the BH procedure given by $\alpha^{*}=[G_{{\tiny\mbox{mixt}}}^{\prime}(0^{+})]^{-1}$; the central limit theorem for the FDP/TDP of BH procedure at level $\alpha$ can only be obtained if $\alpha>\alpha^{*}$ (for $\alpha<\alpha^{*}$, the BH procedure at level $\alpha$ has asymptotically no power). This quantity plays a similar role in the (dependent) conformal setting and we prove results only if $\alpha>\alpha^{*}$.
Finally, as in Neuvial (2008), we consider also the following thresholds
$\mathcal{T}_{\alpha}$ $=\sup\mathopen{}\mathclose{{\left\{t\in(0,1),\ G_{{\tiny\mbox{mixt}}}(t)\geqslant\frac{t}{\alpha}}}\right\};$ (24)
$\mathcal{T}^{w}_{\alpha}$ $=\sup\mathopen{}\mathclose{{\left\{t\in(0,1),\ G_{{\tiny\mbox{mixt}}}^{w}(t)\geqslant\frac{t}{\alpha}}}\right\},$ (25)